A normal distribution of weekly income has a mean of $250 and a standard deviation of $20. Approximately what proportion of people earn between $240 and $260 a week?
Answer by stanbon(57984)

A normal distribution of weekly income has a mean of $250 and a standard deviation of $20. Approximately what proportion of people earn between $240 and $260 a week?
-----
z(240) = (240-250)/20 = -1/2
---
z(260) = (260-250)/20 = +1/2
----
Proportion(240<= x <=260) = P(-1/2 <= z <=1/2) = 0.3829
====================
Cheers,
Stan H.